OpenStudy (anonymous):

Ellen works for a high-speed rail company that wants to develop a new rail line. Ellen’s project is to find a train that is the second fastest in the world. The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. The TGV train in France can reach speeds of 89.44 meters per second. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours. Use complete sentences

OpenStudy (jack1):

1st: convert to same units of measurement as we need the final answer in miles / hr, convert both trains to that UOM. \[\large \frac {320~km} {1~hr} = \frac {???~ miles}{1 hr} \] \[\large \frac {88.44~m} {1~sec} = \frac {???~ miles}{1 hr} \]

OpenStudy (anonymous):

Idk how to do this?

OpenStudy (jack1):

how many miles are there in a kilometer?

OpenStudy (anonymous):

idk

OpenStudy (jack1):

google it then

OpenStudy (anonymous):

1 kilometer=0.621371 miles

OpenStudy (jack1):

cool so to convert, we need to get rid of the kms and replace with miles \[\large \frac {320~km} {1~hr} \times \frac {0.621371~ miles }{1~ km}= \frac {???~ miles}{1 hr} \]

OpenStudy (jack1):

so 320/1 = 320 and 0.621371/1 = 0.621371 so 320 x 0.621371 = answer for train 1

OpenStudy (anonymous):

198.83872 is what i got

OpenStudy (jack1):

cool, so Shinkansen Bullet Train goes at roughly 198.84 miles/hr now the other train how many miles are there in a meter?

OpenStudy (anonymous):

1 meter=0.000621371 miles

OpenStudy (jack1):

perfect dude, so to convert meters to miles, multiply by that number \[\large \frac {88.44~m} {1~sec} \times \frac {0.000621371~miles}{1~m} = \frac {???~ miles}{1 hr} \]

OpenStudy (jack1):

88.44/1 = 88.44 0.000621371/1 = 0.000621371 so 88.44 x 0.000621371 = answer for train 2 in miles per hr

OpenStudy (anonymous):

.05495405124

OpenStudy (jack1):

* sorry man, my bad, that is the answer for miles per second, not miles per hour so to get it to "hours" mulitply .05495405124 x 60 x 60 = answer

OpenStudy (jack1):

\[\large \frac {88.44~m} {1~sec} \times \frac {0.000621371~miles}{1~m} \times \frac {60 sec}{1min} \times \frac {60 min}{1hr} = \frac {???~ miles}{1 hr}\]

OpenStudy (jack1):

so u want the 2nd fastest train so pick something in the middle of these 2 above speeds then work out how far it will travel in 2.5 hrs if its going at the speed you picked